What is the distance between two cars after three hours if one car travels at the rate of 50 kilometers per hour and the other car at 60 km per hour if they have been traveling in the same direction?

Answer by ikleyn(52781)

.

If, in addition, they start from the same common point, then the distance between them in 3 hours is


    3*60 - 3*50 = 3*(60-50) = 3*10 = 30 kilometers.    ANSWER


They separate at the rate of 60 - 50 = 10 kilometers per hour.

Solved, answered and explained.